OpenStudy (anonymous):

The angle of elevation of a ladder leaning against a wall is 72 degree. If the ladder is 20 feet long, how far is the base of the ladder from the wall? A, 6.2 ft b. 16.73 ft C. 19.02 ft D.21.02ft

OpenStudy (anonymous):



OpenStudy (anonymous):

Do you know what to do now that its drawn out?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Do you know that the cosine of an angle equals the adjacent side over the hypotenuse? So: cos(72)=x/20

OpenStudy (anonymous):

i really dont know what to do

OpenStudy (anonymous):

What you have to do is look at triganomitry, you have the length of the hypotanuse and need the length of the side adjacent to the angle - therefore you Cos The angle is 72, so its cos(72) = adj/hyp = adj/20. so times both sides by 20 = 20*cos(72) = the distance of the wall from the base of the ladder Equals 6.18 (2dp)

OpenStudy (anonymous):

Then what